My worksheet says to Determine whether each relation is a function... how do i know what to look for when trying to see if it is a function?
1. y= 1/3x-5
2. y= 2[x]+3 [x] is the absolute value of x, no way to type that in

Answer by CharlesG2(834)

My worksheet says to Determine whether each relation is a function... how do i know what to look for when trying to see if it is a function?
1. y= 1/3x-5
2. y= 2[x]+3 [x] is the absolute value of x, no way to type that in


if a function it will pass the vertical line test, each value of x corresponds to one single calue of y

y = (1/3)x - 5, this is a function, it is a line in form y = mx + b with slope m of 1/3 and y-intercept (vertical-intercept) b of -5

y = 2|x| + 3, used the pipe ("|") symbol for absolute value
x can be positive or negative or zero, but result will always be positive
this is a function, it meets the vertical line test